A computer is purchased for $3,000. In 8 years you scrap it for $o. How much was the computer worth after 5 years? (find equation first)
Solution:

Answer by stanbon(75887)

A computer is purchased for $3,000. In 8 years you scrap it for $o. How much was the computer worth after 5 years? (find equation first)
Solution:
----------------
You have two points: (0,3000) and (8,0)
---
slope = (3000-0)/(0-8) = -375
-----------------------------------
Intercept: (0,3000)
==========================
Equation:
f(x) = -375x+3000
----
f(5) = -375*5+3000 = $1125
=============================
Cheers,
Stan H.